OpenStudy (anonymous):

If h(x) = 3x − 1 and j(x) = −2x, solve h[j(2)] and select the correct answer below. −11 11 −13 13

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

Do you have any idea how to solve?

OpenStudy (anonymous):

ive tried all the ways i thought i knew, but it didnt work, so no :\

OpenStudy (anonymous):

lets start with this what is j(2)=-2(2) j(2)= -4 h[ j(2) ] h[-4] can you solve now?

OpenStudy (anonymous):

-4(-2x(2))= 16x? @Deeezzzz

OpenStudy (danjs):

What is j(2) first?

OpenStudy (danjs):

j(2) = -2*2 = -4 if j(2) = -4 What is F[ j(2)] ...

OpenStudy (danjs):

You start with the inner function, evaluate and work outwards

OpenStudy (danjs):

F[ j(2) ] = F[-4] =

OpenStudy (anonymous):

yea i tried that a few times and could never figure it out

OpenStudy (danjs):

f[-4] = 3(-4) - 1

OpenStudy (danjs):

-12 -1

OpenStudy (danjs):

-13

OpenStudy (danjs):

do you see how to work through these things now?

OpenStudy (anonymous):

oohh yea i see it know

OpenStudy (anonymous):

now

OpenStudy (anonymous):

thankyou

OpenStudy (danjs):

welcdome